### v1

# Energy-Efficient Flying LoRa Gateways: A Multi-Agent Reinforcement Learning Approach

Abstract
With the rapid development of next-generation Internet of Things (NG-IoT) networks, the increasing number of connected devices has led to a surge in power consumption. This rise in energy demand poses significant challenges to resource availability and raises sustainability concerns for large-scale IoT deployments. Efficient energy utilization in communication networks, particularly for power-constrained IoT devices, has thus become a critical area of research. In this paper, we deployed flying LoRa gateways (GWs) mounted on unmanned aerial vehicles (UAVs) to collect data from LoRa end devices (EDs) and transmit it to a central server. Our primary objective is to maximize the global system energy efficiency (EE) of wireless LoRa networks by joint optimization of transmission power (TP), spreading factor (SF), bandwidth (W), and ED association. To solve this challenging problem, we model the problem as a partially observable Markov decision process (POMDP), where each flying LoRa GW acts as a learning agent using a cooperative Multi-Agent Reinforcement Learning (MARL) approach under centralized training and decentralized execution (CTDE). Simulation results demonstrate that our proposed method, based on the multi-agent proximal policy optimization (MAPPO) algorithm, significantly improves the global system EE and surpasses the conventional MARL schemes.
Index Terms:
Internet of Things (IoT), Long rang (LoRa), Energy efficiency, UAV communication, resource allocation.
I Introduction
The next-generation Internet of Things (NG-IoT) technologies for fifth-generation (5G) and sixth-generation (6G) applications are revolutionizing communication by enabling seamless data exchange between devices and networks. This evolution is driving intelligent applications in areas such as healthcare, smart cities, agriculture, and autonomous vehicles [1]. With IoT device connections expected to reach 125 billion by 2030 [2], energy consumption poses a significant challenge, particularly in maintaining low-power and long-range communication networks. Enhancing energy efficiency (EE) is therefore essential to align with global sustainability objectives, including the United Nations Sustainable Development Goal 7: Affordable and Clean Energy [3].
Consequently, low-power wide area networks (LPWANs), particularly long-range (LoRa) technology, have emerged as a key solution to enable energy-efficient communication over long distances. However, because of the extensive interconnectivity, these networks are becoming denser, thus increasing the need for efficient resource allocation techniques. Recent studies on resource allocation approaches, especially in cellular communications, have focused on the use of unmanned aerial vehicles (UAVs) to enhance network coverage and efficiency. Although these approaches have been beneficial in traditional cellular networks, their application to IoT-specific systems has been limited. This limitation has led to inefficiencies, particularly as IoT networks face growing traffic and dynamic communication demands. Addressing these challenges requires the development of efficient resource management algorithms that optimize EE while meeting the dynamic requirements of IoT networks.
A common approach in the literature is alternating optimization (AO), which involves formulating an optimization problem, breaking it down into manageable subproblems, and subsequently developing algorithmic solutions to solve each part [4, 5]. However, this approach becomes particularly critical when considering the complexities introduced by the dense interconnectivity and varying traffic demands of IoT-enabled systems such as flying LoRa gateways (GWs). Moreover, a substantial number of recent studies have used machine learning (ML) techniques to enhance EE in LoRa networks. For example, the authors in [6] address the challenges of LPWAN by using Q-learning with multiple agents framework to dynamically allocate transmission power (TP) and spreading factor (SF) in a LoRa network, thus improving EE. In [7], Q-learning is applied to enhance the packet delivery ratio (PDR) by rewarding successfully received packets. However, reliance on Q-tables makes Q-learning impractical for complex optimization tasks. To address this, proximal policy optimization (PPO) was introduced as a more stable and data-efficient alternative [8]. In [9], the authors propose a DRL-based PPO framework to optimize EE in LoRa networks, utilizing a single GW for data collection to extend the network lifetime.
In this work, we focus on optimizing key resource allocation parameters in LoRa systems, including SF allocation, TP assignment, bandwidth (W) selection, and EDs association, to maximize the global system EE. Our approach incorporates air-to-ground (A2G) link characteristics to account for the impact of the device-to-gateway channel on link quality. To the best of the authors’ knowledge, this is the first work to propose a multi-agent reinforcement learning (MARL) approach that uses multi-agent proximal policy optimization (MAPPO) to enhance global system EE in flying LoRa GWs. The main contributions of this paper are summarized as follows:
  • We formulate the resource allocation and user association problem as a single optimization problem aimed at maximizing EE in LoRa networks, considering both device positions and A2G propagation.
  • We develop a Partially Observable Markov Decision Process (POMDP) model tailored to the LoRa system, with states, action spaces, and reward functions designed for stable MAPPO agent learning.
  • We leverage the MARL-based MAPPO approach that enables flying LoRa GWs to learn a centralized value function while optimizing their policies to maximize global system EE.
  • Our proposed framework leverages one-to-many matching algorithms for an efficient device-to-gateway association, demonstrating faster and more stable convergence compared to state-of-the-art MARL algorithms.
The rest of the paper is organized as follows. We describe the system model in Section II. Section III provides a mathematical formulation of the problem. In Section IV, we describe the proposed MARL configuration in detail. We evaluate the effectiveness of the proposed approach in Section V. Finally, the paper concludes with a comprehensive summary and sets forth some perspectives in Section VI
II System Model
We consider a resource allocation framework for uplink transmission in a LoRa network consisting of $\mathcal{V}$ LoRa EDs, $\mathcal{U}$ flying GWs, and a single network server, as shown in Fig. 1. In this setup, UAVs equipped with LoRa GWs are deployed over a square target area $S$. Each flying LoRa GW has a limited communication range and can simultaneously connect to multiple EDs within its association quota $\Lambda_{max}$. Specifically, each GW collects and decodes packets from all eligible EDs in range and relays these packets to the network server. The sets of EDs and GWs are denoted by $\mathcal{V}=\{1,\ldots,V\}$ and $\mathcal{U}=\{1,\ldots,U\}$, respectively.
II-A EDs Mobility
To accurately model the mobility behavior of our LoRa ground devices, we initially deployed the EDs at random positions within the defined target area $S$. For a given ED $v$, a unique speed vector is assigned $\mathbf{s}_{v}(t)=[s_{v}^{x}(t),s_{v}^{y}(t)]$, where $s_{v}^{x}(t)$ and $s_{v}^{y}(t)$ are respectively the speed along $x$ and $y$ axes. Furthermore, the random assignment ensures that each ED moves in an independent direction with a distinct speed. At each time step $t\in\mathcal{T}=\{1,\ldots,T\}$, the EDs update their positions based on their assigned speed. If an ED reaches the area’s boundary, it bounces back by reversing its speed direction while ensuring it remains within the limits. Additionally, there is a $p_{0}$ probability per ED per time step that its speed will be randomly reassigned, introducing variability in movement over time.
II-B Air-to-Ground Channel Model
In UAV communication systems, A2G link propagation plays a crucial role and can be leveraged for flying GWs communication [10]. In this work, a probabilistic path loss model was used to model A2G communication links, whereby the line-of-sight (LoS) and non-line-of-sight (NLoS) links were considered separately with different probabilities of occurrence. Consequently, the likelihood of having a LoS connection between GW $u$, and an ED $v$ is given by:
$P_{LoS}(\hat{\phi})=\frac{1}{1+\vartheta e^{-\lambda(\hat{\phi}-\vartheta)}},$ (1)
where $\hat{\phi}$ = $sin^{-1}(\frac{h_{u}}{d_{u,v}})$ is the angle of elevation from GWs $u$, $u\in\mathcal{U}$ to EDs $v$, $v\in\mathcal{V}$. $h_{u}$ is the altitude of the UAV $u$. The Euclidean distance between $u$ and $v$ is denoted as $d_{u,v}$. $\vartheta$ and $\lambda$ are constants that depend on the environment. Note, from Eq. (1), the probability values for $P_{NLoS}$ can be expressed as $P_{NLoS}(\hat{\phi})=1-P^{LoS}(\hat{\phi})$. Furthermore, the path loss model from UAV $u$ to the ground LoRa ED $v$ can be expressed as:
$L^{LoS}_{u,v}$ $=L^{FS}(d_{r})+10\delta_{LoS}d_{u,v}+\mathcal{X}\sigma_{LoS},$ (2a)
$L^{NLoS}_{u,v}$ $=L^{FS}(d_{r})+10\delta_{NLoS}d_{u,v}+\mathcal{X}\sigma_{NLoS},$ (2b)
where $L^{FS}$ denotes the free-space path-loss with $L^{FS}(d_{r})=20\log_{10}\left(\frac{4\pi d_{r}f}{c}\right)$. $\delta$ is the path loss exponent. $f$ is the carrier frequency in Hz, $c$ in m/s is the speed of light, and $d_{r}$ is the reference distance. $\mathcal{X}\sigma_{LoS}$ and $\mathcal{X}\sigma_{NLoS}$ are the shadowing random variables which are characterized as Gaussian random variables with zero mean and $\sigma_{LoS}$ and $\sigma_{NLoS}$ standard deviations. Consequently, the overall A2G path loss between GW $u$ and LoRa ED $v$ is characterized as:
$l_{u,v}^{a2g}=P_{LoS}(\hat{\phi})L^{LoS}_{u,v}+P_{NLoS}(\hat{\phi})L^{NLoS}_{u,v}.$ (3)
II-C Energy Efficiency Model
To model the EE of our proposed system, we begin by linking the network topology to the LoRa PHY parameters. This is achieved by first modeling the signal-to-noise ratio (SNR) between UAV $u$ and ED $v$ at time slot $t$, expressed as
$\rho_{u,v}(t)=\frac{P_{v}(t)\cdot G_{u,v}}{\sigma^{2}},$ (4)
where $G_{u,v}=10^{-{l_{u,v}^{\text{a2g}}}/{10}}$ represents the channel gain, $P_{v}(t)$ is the TP of an ED at time $t$, and $\sigma^{2}$ is the noise power. Building on this, we calculate the signal-to-interference-plus-noise ratio (SINR) $\mho_{u,v}^{n}(t)$ between UAV $u$ and ED $v$ using the $n$-th SF at time slot $t$ as follows:
$\mho_{u,v}^{n}(t)=\frac{\rho_{u,v}(t)}{\sum_{{v}^{\prime}\in\mathcal{V}\setminus\{v\}}\psi_{{v}^{\prime},n}(t)\cdot\rho_{a_{{v}^{\prime}}(t),{v}^{\prime}}(t)+1},$ (5)
where $\rho_{a_{{v}^{\prime}}(t),{v}^{\prime}}(t)$ represents the SNR between ED $v^{\prime}$ and its associated UAV $a_{{v}^{\prime}}(t)$, and $\psi_{{v}^{\prime},n}(t)$ is the binary association parameter that indicates whether user $v^{\prime}$ selected SF $n$.
Figure 1: The studied system model.
The achievable data rate for the link between UAV $u$ and ED $v$ at time slot $t$ is derived using the Shannon-Hartley theorem [11] and is given by:
$\Re_{u,v}(t)=W_{v}(t)\cdot\log_{2}\left(1+\mho_{u,v}^{n}(t)\right),$ (6)
where $W_{v}(t)$ is the bandwidth allocated to the communication link.
Furthermore, the EE of each UAV is calculated by dividing the sum of all uplink data rates from all EDs connected to the UAV by the total consumed power. Hence, we define the EE $\zeta_{u}(t)$ of UAV $u$ at time $t$ as:
$\zeta_{u}(t)=\frac{\sum_{v\in\mathcal{V}}\Re_{u,v}(t)a_{u,v}(t)}{P_{T}+P_{c}},$ (7)
where $\Re$ in bits/second is the data rate given in Eq.(6). $P_{c}$ is the circuit power consumption, and $P_{T}=\sum_{v=1}^{V}P_{v}(t)a_{u,v}(t)$ is the transmission power, where $a_{u,v}(t)$ is the ED binary association.
II-D EDs Association and Resource Allocation Scheme
In the UAV-mounted LoRa GWs system, implementing a dynamic EDs association is a crucial element in efficiently managing connections between ground EDs and flying GWs. Given that each UAV serves multiple EDs, it is important to ensure the association is properly established to ensure load balance in the network. Therefore, the binary association between a UAV $u$ and an ED $v$ at time slot $t$ is denoted by $a_{u,v}(t)\in\{0,1\}$, $a_{u,v}(t)=1$ if ED $v$ is being served by UAV $u$ at time $t$, and $a_{u,v}(t)=0$ otherwise. The index of the UAV selected by ED $v$ at time $t$ can be expressed as $a_{v}(t)=\sum_{u\in\mathcal{U}}a_{u,v}(t)\cdot u.$ [^1] This scalar value represents the UAV associated with ED $v$ and is essential for calculating interference in the SINR as given in Eq. (5).
In this work, SF is selected from a vector $\mathbf{\Psi}=\{\psi_{1},\cdots,\psi_{N}\}$. For each ED $v$, the allocation of SF follows a binary association that is expressed as $\psi_{v,n}(t)$, $n\in\mathcal{N}=\{1,\cdots,N\}$. Therefore, $\psi_{v,n}(t)=1$ if ED $v$ communicates at SF $\psi_{n}$ during time slot $t$; otherwise, $\psi_{v,n}(t)=0$. The selected index can be expressed as $\Psi_{v}(t)=\sum_{n\in\mathcal{N}}\psi_{v,n}(t)\cdot\psi_{n}.$ We assume that each ED can transmit data using only one SF at any given time step, leading to the following constraint
$\sum_{n=1}^{N}\psi_{v,n}(t)\leq 1,\forall v\in\mathcal{V}.$ (8)
Similarly, the transmission power level is selected from a vector $\mathbf{P}=\{p_{1},\cdots,p_{J}\}$ in dBm. For each ED $v$, we define a binary allocation variable $p_{v,j}(t)$, $j\in\mathcal{J}=\{1,\cdots,J\}$. Therefore, $p_{v,j}(t)=1$ if ED $v$ transmits with power level $p_{j}$ at time $t$; and $p_{v,j}(t)=0$ otherwise. Note that the selected index is $P_{v}(t)=\sum_{j\in\mathcal{J}}p_{v,j}(t)\cdot p_{j}.$ We also assume that each ED can transmit data using only one TP at each time step, imposing the following constraints:
$\sum_{j=1}^{\mathcal{J}}p_{v,j}(t)\leq 1,\forall v\in\mathcal{V}.$ (9)
In addition, the bandwidth $W$ for communication is selected from a vector $\mathbf{W}=\{w_{1},\cdots,w_{M}\}$ in kHz. We assume that each deployed flying GW contains a distinct LoRa module with specific bandwidth requirements. Hence, the bandwidth binary allocation is $w_{v,m}(t)$, $m\in\mathcal{M}=\{1,\cdots,M\}$ such that $w_{v,m}(t)=1$ if ED $v$ transmits using bandwidth $w_{m}$ at time $t$; otherwise, $w_{v,m}(t)=0$. Also, the selected index used in Eq.(6) is expressed as $W_{v}(t)=\sum_{m=1}^{M}w_{v,m}(t)\cdot w_{m}.$
Consequently, we define finite sets for all possible SF selections $\bar{\bf\Psi}$, TP allocations $\bar{\textbf{P}}$, user associations a, and bandwidth selections $\bar{\textbf{W}}$, which can be expressed as:
$\bar{\mathbf{\Psi}}=\{\Psi_{v}(t)\in{\bf\Psi}\mid\sum_{n=1}^{\mathcal{N}}\psi_{v,n}(t)\leq 1,\quad\forall v\in\mathcal{V}\},$ (10a)
$\bar{\mathbf{P}}=\{P_{v}(t)\in\textbf{P}|\sum_{j=1}^{\mathcal{J}}p_{v,j}(t)\leq 1,\quad\forall v\in\mathcal{V}\},$ (10b)
$\mathbf{a}=\{a_{v}(t)\in\mathcal{U}|\sum_{u=1}^{\mathcal{U}}a_{u,v}(t)\leq 1,\quad\forall v\in\mathcal{V}\},$ (10c)
$\bar{\mathbf{W}}=\{W_{v}(t)\in\textbf{W}|\sum_{m=1}^{\mathcal{M}}w_{v,m}(t)\leq 1,\quad\forall v\in\mathcal{V}\}.$ (10d)
III Problem Formulation
This section presents a detailed description of the optimization problem based on our proposed system model. The objective of the paper is to maximize the global system EE in flying LoRa GWs, as defined in Eq.(11a). Consequently, our optimization problem is formulated as follows.
[…]
$\;a_{u,v}(t)\in\{0,1\},\forall u,v,$ (11c)
$\;\sum_{u\in\mathcal{U}}a_{u,v}(t)\leq 1,\forall v,$ (11d)
$\;\sum_{v\in\mathcal{V}}a_{u,v}(t)\leq\Lambda_{max},\forall u,$ (11e)
$\;\psi_{v}(t)\in{\bf\Psi},\forall v,$ (11f)
$\;\mho_{u,v}\geq\bar{\mho}_{SF},\forall u,v,$ (11g)
$\;w_{v,m}(t)\in\textbf{W},\forall v,$ (11h)
[…]
Constraint  (11b) limits the lower and the upper bounds of the transmission power available for each LoRa ED. The binary indicator of ED $a_{u,v}$ in  (11c), is either 0 or 1. In constraint  (11d), each ED can access at most one channel. Constraint  (11e) ensures that at most, there are $\Lambda_{max}$ EDs $v$ that can be associated to a single gateway $u$. Furthermore, constraint  (11f) specifies the available set of SFs. It is important to note that constraint  (11g) ensures that the received SINR must not be smaller than the threshold $\bar{\mho}$ shown in Table I, to detect the required LoRa EDs adopting a particular SF correctly [6] and finally, constraint  (11h) ensures that the bandwidth selected for each ED $v$ is from the discrete set W.
TABLE I: SINR Thresholds ¯$\bar{\mho}_{SF}$ with $W=125$ kHz [6]
SF 7 8 9 10 11 12
¯$\bar{\mho}_{SF}$ (dB) -7.5 -10 -12.5 -15 -17.5 -20
The formulated optimization problem (P1) is NP-hard due to its combination of binary and stochastic constraints, along with the selection of discrete variables for SF, TP, and W, leading to a non-convex objective function that is challenging for conventional optimization methods. As established in [12], no efficient polynomial-time solution exists for this problem. Consequently, traditional RL methods, which require storing all MDP tuples in a table, are computationally intensive and unsuitable for multi-agent scenarios; thus, we design POMDP method and leverage MAPPO-based algorithm to tackle the formulated problem in Eq.(11a).
IV Multi-Agent PPO Algorithm
MAPPO is a promising framework that builds on the centralized training, decentralized execution (CTDE) concept and extends the PPO to the multi-agent system [13, 14]. It focuses on the actor-critic architecture, which consists of two components: the actor, who is in charge of making decisions, and the critic, who analyses those actions using a value function. In a CTDE architecture shown in Fig. 2, the critic network learns a centralized value function that includes knowledge of all agents’ activities, whilst each agent uses the actor-network to determine its policy based solely on local observations. Furthermore, due to the limited communication range and the maximum number of EDs each flying LoRa GW can associate with at any given time, the global state of the system cannot be fully observed by a single agent. Consequently, we model the problem using a partially observable Markov decision process (POMDP). In this framework, each agent only has partial information about the environment, leading to uncertainty in decision-making. Formally, the POMDP is defined by the tuple $\left\langle\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma,\mathcal{O}\right\rangle$.
Hence, the proposed framework consists of the following elements:
1) Agents $\mathcal{U}$: The set of flying LoRa GWs.
2) States $\mathcal{S}$: The global state set $\mathcal{S}(t)$ represents the environment’s current conditions at time step $t$. The global state is given by $\mathcal{S}(t)=\{\Psi,P,W,\Re,assEDs_{pos},GWs_{pos}\}$. These elements correspond to the assigned SF, TP, bandwidth, the computed data rate, the 2D positions of the associated EDs, and the positions of the GWs.
3) Observations $\mathcal{O}$: Due to partial observability, each flying LoRa GW $u$ at time step $t$ perceives only a subset of the global state $\mathcal{S}(t)$. Therefore, the observation space $o_{u}(t)\in\mathcal{S}$ for agent $u$ is defined as $o_{u}(t)=\{\Psi_{v}(t),P_{v}(t),\Re_{v}(t),assEDs_{pos,v}(t),GWs_{neighbors}(t)\}$. Here, the set $assEDs_{pos,v}$ represents the positions of the EDs currently associated with GW $u$. This partial observation mechanism ensures that each GW makes decisions based on local information, requiring cooperative strategies for optimal network-wide performance.
4) Actions $\mathcal{A}$: At each time step $t$, each agent makes decisions regarding resource allocation. These actions include selecting the appropriate SF $\psi(t)$, TP $p(t)$, and bandwidth $w(t)$. The actions taken by an agent can be expressed as $A_{u}(t)=(\psi(t),p(t),w(t))$. Once a decision has been made, the selected LoRa ED is served using the allocated resources.
5) Rewards $\mathcal{R}$: The reward function is related to the global EE system of all flying LoRa GWs, as described in 11a. There is defined as:
$r_{u}(t)=\sum_{t=1}^{T}\sum_{u=1}^{U}\frac{\sum_{v\in\mathcal{V}}\Re_{u,v}(t)a_{u,v}(t)}{\sum_{v\in\mathcal{V}}P_{u,v}(t)a_{u,v}(t)+P_{c}}.$ (12)
We ensure that the constraints in (11f) and (11g) are respected before rewards are assigned to the agents. Specifically, when the observation ${o}_{u}(t)$ is received, the agent must take an action ${a}_{t}\in\mathcal{A}$ that satisfies the constraints in P1.
5) Policy function: The policy function $\boldsymbol{\pi_{\theta_{u}}}(\textbf{a}_{u},o_{u})$ is modeled by an actor-network, with the vector $\boldsymbol{\theta}_{u}$ as its parameters. It determines the strategy for the flying GWs based on their local observations.
6) Value function: The value function $V_{\phi}(\boldsymbol{o_{u}})$ is implemented by a critic network parameterized by $\phi$. This network evaluates the expected future rewards for flying GWs given the current state $\boldsymbol{o}_{u}$. The critic aims to learn a value function that approximates the optimal future rewards, guiding the agents toward the globally optimal policy that maximizes their long-term rewards.
Figure 2: Illustration of Centralized Critic and Decentralized Actor for MAPPO training.
As shown in Fig. 2, our system leverages MAPPO, which is based on the CTDE framework. During the training phase, MAPPO alternates between optimizing the actor and critic networks until stable convergence is achieved. Specifically, the flying GWs locally update the policy actor using the PPO method. At each training step, given the state of a GW, the actor-network is trained, and action is sampled according to the policy function $\pi_{\theta_{u}}(a_{u}(t),o_{u}(t))$. Once the joint action is executed, the corresponding reward is observed. Subsequently, the global state vector, representing the collective states of all flying GWs, is passed to the critic network. The critic network is trained by minimizing a predefined loss function, which helps evaluate future rewards for the policy and improves the overall strategy over time.
V Performance Evaluation
V-A Simulation setup
[…]
In our simulation, we compare our proposed approach with three other MARL algorithms: Counterfactual Multi-Agent (COMA) [15], Multi-Agent Advantage Actor-Critic (MAA2C) [16], and Value-Decomposition Networks (VDN)[17]. The COMA algorithm is an actor-critic method in which a centralized critic estimates the Q-function, with decentralized actors optimizing their policies. The MAA2C algorithm extends the A2C algorithm to multi-agent scenarios by incorporating a centralized critic that learns the joint value function. At the same time, each agent maintains its own actor to learn individual policies. Finally, the VDN framework is designed for cooperative MARL tasks, wherein each agent learns a distinct value function that is decomposed into shared and local value functions.
V-B Simulation Results
In Fig. 3(a), we illustrate the association between EDs and GWs in a 3D plane. As shown in the figure, we employ the one-to-many matching scheme from [18], where each GW is associated with multiple LoRa EDs. Consequently, the quota constraint is maintained, and the network load is balanced across GWs. In Fig. 3(b), the training performance is evaluated for four different learning rates ($\alpha$). In this setup, we fixed the clipping range at $\epsilon=0.2$, with 60 EDs and 5 GWs. Each training curve in the figure represents the global cumulative reward over environment timesteps for a specific learning rate: $\alpha=0.0005$, $\alpha=0.0003$, $\alpha=0.003$, and $\alpha=0.03$. It can be observed that $\alpha=0.0003$ achieves the highest rewards with stable convergence throughout the training phase. In contrast, $\alpha=0.03$ performs the worst, yielding very low global cumulative rewards with an unstable training process at different timesteps. Although $\alpha=0.0005$ converges slightly faster than $\alpha=0.0003$, it becomes unstable toward the end of training. Therefore, we utilize $\alpha=0.0003$ as it achieves the highest reward while maintaining stable learning.
In Fig. 3(c), the convergence behavior is examined for different numbers of GWs while keeping EDs fixed at 60, $\epsilon=0.2$, and $\alpha=0.0003$. The results show that with fewer GWs (e.g., 2), the rewards converge faster and stabilize earlier but fail to achieve the highest global system EE. On the other hand, as the number of GWs increases (e.g., 6 and 8), the convergence rate decreases. However, it can be observed that a higher number of GWs leads to greater rewards. This is expected, as more GWs provide better spatial coverage and improved connectivity, resulting in more efficient data aggregation and reduced communication overhead. Consequently, this enhances the overall global system EE and enables better resource allocation.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 3: (a) 3D plane of LoRa ED-gateway association, (b) Training reward with different learning rates $\alpha$, (c) Training reward with different gateways.
In Fig. 4, we compare the optimal EE versus the number of active EDs across multiple multi-agent RL benchmarks, including COMA, MAA2C, and VDN. Our proposed framework consistently achieves the highest global EE due to its CTDE paradigm, which enables efficient EE allocation across the network. Unlike COMA, which suffers from high variance and scalability issues, MAPPO ensures balanced reward distribution. On the other hand, MAA2C outperforms COMA and VDN, its cooperative learning limitations result in declining EE with an increasing number of EDs. Similarly, VDN exhibits the lowest EE due to its lack of explicit agent coordination. Additionally, with five fixed GWs, our approach maintains superior global system EE, whereas other algorithms struggle in denser networks due to coordination failures and communication overhead. It is important to note that as the number of EDs increases, the global EE naturally decreases. A higher number of EDs leads to a higher total TP consumption, increased interference, and reduced data rates, all of which contribute to lower EE. Despite these challenges, MAPPO mitigates this degradation more effectively than the other approaches. In comparison with the closest-performing algorithm (MAA2C), our approach improves the global system EE by 30.5%, 7.2%, 6.5%, 10.9%, and 11.0% for 20, 30, 40, 50, and 60 EDs, respectively.
Figure 4: Total energy efficiency under different number of end-devices
VI Conclusion
In this paper, we considered the joint optimization of assigning SF, TP, BW, and ED associations between multiple EDs and GWs while considering resource constraints and the A2G wireless link to optimize the global system EE in a multi-flying LoRa network. We showed that the resulting sequential decision-making problem can be modeled as a POMDP, and we proposed a model-free RL algorithm that leverages a novel MAPPO scheme. We used simulation to show that our proposed approach can learn a good policy approximation by optimizing the global system EE in LoRa networks. Interesting directions of future work include integrating trajectory optimization for multiple flying LoRa gateways for efficient resource management. Our approach could be extended by exploring the Age of Information (AoI) scheme to enhance the freshness of data collection.
VII Acknowledgement
This work was sponsored by the Junior Faculty Development program under the UM6P-EPFL Excellence in Africa Initiative.