### v3

# Energy-Efficient Flying LoRa Gateways: A Multi-Agent Reinforcement Learning Approach

Abstract
As next-generation Internet of Things (NG-IoT) networks continue to grow, the number of connected devices is rapidly increasing, along with their energy demands. This creates challenges for resource management and sustainability. Energy-efficient communication, particularly for power-limited IoT devices, is therefore a key research focus. In this paper, we deployed flying LoRa gateways mounted on unmanned aerial vehicles (UAVs) to collect data from LoRa end devices and transmit it to a central server. Our primary objective is to maximize the global system energy efficiency of wireless LoRa networks by joint optimization of transmission power, spreading factor, bandwidth, and user association. To solve this challenging problem, we model the problem as a partially observable Markov decision process (POMDP), where each flying LoRa GW acts as a learning agent using a cooperative multi-agent reinforcement learning (MARL). Simulation results demonstrate that our proposed method, based on the multi-agent proximal policy optimization algorithm, significantly improves the global system energy efficiency and surpasses the popular MARL and other conventional schemes.
Index Terms:
Internet of Things (IoT), Long Rang (LoRa), Energy efficiency, UAV, Resource allocation.
I Introduction
The next-generation Internet of Things (NG-IoT) technologies for 5G and 6G applications are revolutionizing communication by enabling seamless data exchange between devices and networks. This evolution drives intelligent applications in areas such as healthcare, smart cities, agriculture, and autonomous vehicles [1]. With IoT device connections expected to reach 125 billion by 2030 [2], energy consumption presents a significant challenge, particularly in maintaining low-power and long-range communication networks. Enhancing energy efficiency is therefore essential to align with global sustainability objectives, including the United Nations Sustainable Development Goal 7: Affordable and Clean Energy [3].
Low-power wide area networks (LPWANs), particularly long-range (LoRa) technology, have emerged as a cost-effective solution for long-distance communication in IoT applications. However, existing terrestrial LoRa networks depend on fixed ground-based gateways (GWs), which struggle with non-line-of-sight (NLoS) propagation, especially in dense urban or remote environments. While deploying additional terrestrial LoRa GWs is affordable, it does not necessarily resolve NLoS issues. Conversely, satellite-based IoT solutions, such as the FOSSA system [^1] FOSSA systems is a low-Earth orbit (LEO) satellite network providing global IoT connectivity for remote areas. More details at: https://fossa.systems/, aim at connecting IoT devices to non-terrestrial networks. However, this approach significantly increases transmission power requirements and introduces higher latency, making it impractical for many energy-constrained IoT applications.
Beyond infrastructure deployment, effective resource allocation plays a crucial role in optimizing LoRa network performance. Existing studies primarily rely on alternative optimization techniques, where complex optimization problems are decomposed into sub-problems and solved iteratively [4]. Although alternative optimization methods have proven effective in certain network environments, they often struggle to adapt dynamically to varying IoT traffic patterns and environmental conditions, leading to suboptimal resource utilization. For example, the authors in [5] employ an alternative optimization-based method for a single-flying LoRa GW, which lacks adaptability in dynamic environments. In contrast, reinforcement learning (RL)-based approaches, such as those in [6, 7], utilize the Q-learning approach for resource allocation. However, these methods depend on static Q-tables, making them impractical for managing complex and dynamic IoT environments. To improve the energy efficiency system of LoRa, the work in [8] proposes a framework based on deep reinforcement learning (DRL) proximal policy optimization (PPO), but it is limited to a single GW for data collection, which restricts scalability and flexibility.
To address these limitations, we propose a UAV-assisted multi-flying LoRa GW deployment. This novel approach can dynamically reposition GWs to enhance network coverage, mitigate NLoS issues, optimize end-devices (EDs) association, minimize transmission power, and maximize energy-efficient communication. Furthermore, we utilize a multi-agent reinforcement learning (MARL) framework with multi-agent PPO (MAPPO) to optimize resource allocation in multi-flying LoRa GWs. Although single-agent RL approaches have shown promise in simpler scenarios [8], extending them to multi-UAV LoRa networks introduces significant challenges, including partial observability, decentralized control, and the need to optimize multiple interdependent parameters in dynamic environments. Our MAPPO approach overcomes these challenges by allowing UAV-mounted GWs to autonomously manage key network parameters, including spreading factor (SF) allocation, transmission power (TP) control, bandwidth (W) selection, and ED association. By incorporating air-to-ground (A2G) link characteristics, our framework ensures efficient resource distribution while maximizing system-wide energy efficiency. Unlike traditional methods, MAPPO continuously updates policies in real time, making it more adaptable to dynamic network conditions. To the best of the authors’ knowledge, this is the first work to propose a MARL-based approach for optimizing global system energy efficiency in a multi-flying LoRa GW deployment. The main contributions of this paper are summarized as follows:
  • We formulate the resource allocation and user association problem as a single optimization problem aimed at maximizing energy efficiency in LoRa networks, considering both device positions and A2G propagation.
  • We develop a partially observable Markov decision process (POMDP) model tailored to the LoRa system, with states, action spaces, and reward functions designed for stable MAPPO agent learning.
  • We leverage the MARL-based MAPPO approach that enables flying LoRa GWs to learn a centralized value function while optimizing their policies to maximize global system energy efficiency.
  • Our proposed framework leverages one-to-many matching algorithms for an efficient device-to-gateway association, demonstrating faster and more stable convergence compared to state-of-the-art MARL algorithms and other conventional allocation schemes.
The rest of the paper is organized as follows. We describe the system model in Section II. Section III provides a mathematical formulation of the problem. In Section IV, we describe the proposed MARL configuration in detail. We evaluate the effectiveness of the proposed approach in Section V. Finally, the paper concludes with a comprehensive summary and sets forth some perspectives in Section VI
II System Model
We consider a resource allocation framework for uplink transmission in a LoRa network consisting of $\mathcal{V}$ LoRa EDs, $\mathcal{U}$ flying GWs, and a single network server, as shown in Fig. 1. In this setup, UAVs equipped with LoRa GWs are deployed over a square target area $S$. Each flying LoRa GW has a limited communication range and can simultaneously connect to multiple EDs within its association quota $\Lambda_{max}$. Specifically, each GW collects and decodes packets from all associated EDs in range and relays these packets to the network server. The sets of EDs and GWs are denoted by $\mathcal{V}=\{1,\ldots,V\}$ and $\mathcal{U}=\{1,\ldots,U\}$, respectively.
II-A LoRa End Devices Mobility
To accurately model the mobility behavior of our LoRa ground EDs, we initially deployed the EDs at random positions within an area of interest. For a given ED $v$, a unique speed vector is assigned $\mathbf{s}_{v}(t)=[s_{v}^{x}(t),s_{v}^{y}(t)]$, where $s_{v}^{x}(t)$ and $s_{v}^{y}(t)$ are respectively the speed along $x$ and $y$ axes. Furthermore, the random assignment ensures that each ED moves in an independent direction with a distinct speed. At each time step $t\in\mathcal{T}=\{1,\ldots,T\}$, the EDs update their positions based on their assigned speed. If an ED reaches the area’s boundary, it bounces back by reversing its speed direction while ensuring it remains within the limits. Additionally, there is a $p_{0}$ probability per ED per time step that its speed will be randomly reassigned, introducing variability in movement over time.
Figure 1: The studied system model.
II-B Air-to-Ground Channel Model
In this study, we adopt a probabilistic path-loss model for A2G links, accounting separately for Line-of-Sight (LoS) and NLoS propagation scenarios based on their occurrence probabilities. Consequently, the probability of having a LoS connection between GW $u\in\mathcal{U}$ and an ED $v\in\mathcal{V}$ is computed as:
$P_{\text{LoS}}(\hat{\phi})=\frac{1}{1+\alpha e^{-\lambda(\hat{\phi}-\alpha)}},$ (1)
where $\hat{\phi}=\tan^{-1}\left(\frac{h_{u}}{d_{u,v}}\right)$ represents the elevation angle between UAV $u$ at altitude $h_{u}$ and ED $v$. The horizontal Euclidean distance $d_{u,v}$ between UAV GW $u$ and ED $v$ is calculated based on their Cartesian coordinates as $d_{u,v}=\sqrt{(x_{u}-x_{v})^{2}+(y_{u}-y_{v})^{2}},$ where $(x_{u},y_{u})$ and $(x_{v},y_{v})$ represent the horizontal positions of UAV $u$ and ED $v$, respectively. Here, the constants $\alpha$ and $\lambda$ are determined empirically based on environmental conditions. Note, from Eq.(1), the probability of the NLoS can be expressed as $P_{\text{NLoS}}(\hat{\phi})=1-P_{\text{LoS}}(\hat{\phi})$.
Furthermore, we consider a free-space path loss (FSPL) at distance $d_{u,v}$ with carrier frequency $f$ as $L^{\text{FS}}(d_{u,v})=20\log_{10}\left(\frac{4\pi d_{u,v}f}{c}\right).$ Here, $c$ is the speed of light. Finally, we model the overall A2G path loss according to [9], by incorporating both LoS and NLoS conditions given as:
$L_{u,v}^{\text{A2G}}=L^{\text{FS}}(d_{u,v})+\eta_{\text{LoS}}P_{\text{LoS}}(\hat{\phi})+\eta_{\text{NLoS}}P_{\text{NLoS}}(\hat{\phi}),$ (2)
where $\eta_{\text{LoS}}$ and $\eta_{\text{NLoS}}$ are additional loss factors associated with LoS and NLoS conditions, respectively.
II-C Energy Efficiency Model
To model the energy efficiency of our proposed system, we begin by linking the network topology to the LoRa PHY parameters. This is achieved by first modeling the signal-to-noise ratio (SNR) between UAV $u$ and ED $v$ at time slot $t$, expressed as $\rho_{u,v}(t)=\frac{P_{v}(t)\cdot G_{u,v}}{\sigma^{2}}$, where $G_{u,v}=10^{-{L_{u,v}^{\text{A2G}}}/{10}}$ represents the channel gain, $P_{v}(t)$ is the TP of an ED at time $t$, and $\sigma^{2}$ is the noise power. Building on this, we calculate the signal-to-interference-plus-noise ratio (SINR) $\mho_{u,v}^{n}(t)$ between UAV $u$ and ED $v$ using the $n$-th SF at time slot $t$ as follows:
$\mho_{u,v}^{n}(t)=\frac{\rho_{u,v}(t)}{\sum_{{v}^{\prime}\in\mathcal{V}\setminus\{v\}}\psi_{{v}^{\prime},n}(t)\cdot\rho_{a_{{v}^{\prime}}(t),{v}^{\prime}}(t)+1},$ (3)
where $\rho_{a_{{v}^{\prime}}(t),{v}^{\prime}}(t)$ represents the SNR between ED $v^{\prime}$ and its associated UAV $a_{{v}^{\prime}}(t)$, and $\psi_{{v}^{\prime},n}(t)$ is the binary association parameter that indicates whether user $v^{\prime}$ selected SF $n$.
The achievable data rate for the link between UAV $u$ and ED $v$ at time slot $t$ is derived using the Shannon-Hartley theorem [10] and is given by:
$\Re_{u,v}(t)=W_{v}(t)\cdot\log_{2}\left(1+\mho_{u,v}^{n}(t)\right),$ (4)
where $W_{v}(t)$ is the bandwidth allocated to the communication link. Furthermore, the energy efficiency of each UAV is calculated by dividing the sum of all uplink data rates from all EDs connected to the UAV by the total consumed power. Hence, we define the energy efficiency $\zeta_{u}(t)$ of UAV $u$ at time $t$ as:
$\zeta_{u}(t)=\frac{\sum_{v\in\mathcal{V}}\Re_{u,v}(t)a_{u,v}(t)}{P_{T}+P_{c}},$ (5)
where $\Re$ in bits/second is the data rate given in Eq.(4). $P_{c}$ is the circuit power consumption, and $P_{T}=\sum_{v=1}^{V}P_{v}(t)a_{u,v}(t)$ is the transmission power, where $a_{u,v}(t)$ is the ED binary association.
II-D EDs Association and Resource Allocation Scheme
In the UAV-mounted LoRa GWs system, implementing a dynamic EDs association is a crucial element in efficiently managing connections between ground EDs and flying GWs. Given that each UAV serves multiple EDs, it is important to ensure the association is properly established to ensure load balance in the network. Therefore, the binary association between a UAV $u$ and an ED $v$ at time slot $t$ is denoted by $a_{u,v}(t)\in\{0,1\}$, $a_{u,v}(t)=1$ if ED $v$ is being served by UAV $u$ at time $t$, and $a_{u,v}(t)=0$ otherwise. The index of the UAV selected by ED $v$ at time $t$ can be expressed as $a_{v}(t)=\sum_{u\in\mathcal{U}}a_{u,v}(t)\cdot u$.
In this work, SF is selected from a vector $\mathbf{\Psi}=\{\psi_{1},\cdots,\psi_{N}\}$. For each ED $v$, the allocation of SF follows a binary association that is expressed as $\psi_{v,n}(t)$, $n\in\mathcal{N}=\{1,\cdots,N\}$. Therefore, $\psi_{v,n}(t)=1$ if ED $v$ communicates at SF $\psi_{n}$ during time slot $t$; otherwise, $\psi_{v,n}(t)=0$. The selected index can be expressed as $\Psi_{v}(t)=\sum_{n\in\mathcal{N}}\psi_{v,n}(t)\cdot\psi_{n}.$ We assume that each ED can transmit data using only one SF at any given time step, leading to the following constraint
$\sum_{n=1}^{N}\psi_{v,n}(t)\leq 1,\forall v\in\mathcal{V},t\in\mathcal{T}.$ (6)
Similarly, the transmission power level is selected from a vector $\mathbf{P}=\{p_{1},\cdots,p_{J}\}$ in dBm. For each ED $v$, we define a binary allocation variable $p_{v,j}(t)$, $j\in\mathcal{J}=\{1,\cdots,J\}$. Therefore, $p_{v,j}(t)=1$ if ED $v$ transmits with power level $p_{j}$ at time $t$; and $p_{v,j}(t)=0$ otherwise. Note that the selected index is $P_{v}(t)=\sum_{j\in\mathcal{J}}p_{v,j}(t)\cdot p_{j}.$ We also assume that each ED can transmit data using only one TP at each time step, imposing the following constraints:
$\sum_{j=1}^{\mathcal{J}}p_{v,j}(t)\leq 1,\forall v\in\mathcal{V},t\in\mathcal{T}.$ (7)
In addition, the bandwidth $W$ for communication is selected from a vector $\mathbf{W}=\{w_{1},\cdots,w_{M}\}$ in kHz. We assume that each deployed flying GW contains a distinct LoRa module with specific bandwidth requirements. Hence, the bandwidth binary allocation is $w_{v,m}(t)$, $m\in\mathcal{M}=\{1,\cdots,M\}$ such that $w_{v,m}(t)=1$ if ED $v$ transmits using bandwidth $w_{m}$ at time $t$; otherwise, $w_{v,m}(t)=0$. Also, the selected index used in Eq.(4) is expressed as $W_{v}(t)=\sum_{m=1}^{M}w_{v,m}(t)\cdot w_{m}.$
Consequently, we define finite sets for all possible SF selections $\bar{\bf\Psi}$, TP allocations $\bar{\textbf{P}}$, user associations a, and bandwidth selections $\bar{\textbf{W}}$, which can be expressed as:
$\bar{\mathbf{\Psi}}=\{\Psi_{v}(t)\in{\bf\Psi}\mid\sum_{n=1}^{\mathcal{N}}\psi_{v,n}(t)\leq 1,\quad\forall v\in\mathcal{V}\},$ (8a)
$\bar{\mathbf{P}}=\{P_{v}(t)\in\textbf{P}|\sum_{j=1}^{\mathcal{J}}p_{v,j}(t)\leq 1,\quad\forall v\in\mathcal{V}\},$ (8b)
$\mathbf{a}=\{a_{v}(t)\in\mathcal{U}|\sum_{u=1}^{\mathcal{U}}a_{u,v}(t)\leq 1,\quad\forall v\in\mathcal{V}\},$ (8c)
$\bar{\mathbf{W}}=\{W_{v}(t)\in\textbf{W}|\sum_{m=1}^{\mathcal{M}}w_{v,m}(t)\leq 1,\quad\forall v\in\mathcal{V}\}.$ (8d)
III Problem Formulation
This section provides a detailed description of the optimization problem based on our proposed system model. The objective of this paper is to maximize the global system energy efficiency of flying LoRa GWs, as defined in Eq. (9a). Consequently, we formulate our optimization problem as follows:
$\max_{\mathbf{\bar{\Psi},\bar{P},a,\bar{W}}}$ $\quad\sum_{t=1}^{T}\sum_{u=1}^{U}\zeta_{u}(t),$ (9a)
s.t. $\;a_{u,v}(t)\in\{0,1\},\forall u,v,t\in\mathcal{T},$ (9b)
$\;\sum_{u\in\mathcal{U}}a_{u,v}(t)\leq 1,\forall v,t\in\mathcal{T},$ (9c)
$\;\sum_{v\in\mathcal{V}}a_{u,v}(t)\leq\Lambda_{max},\forall u,t\in\mathcal{T},$ (9d)
$\;\psi_{v,n}(t)\in{\bf\Psi},\forall v,t\in\mathcal{T},$ (9e)
$\;p_{v,j}\in\textbf{P},\forall v,$ (9f)
$\;w_{v,m}(t)\in\textbf{W},\forall v,t\in\mathcal{T},$ (9g)
$\;\rho_{u,v}(t)\geq\text{SNR}_{threshold},\forall u,v,t\in\mathcal{T}$ (9h)
[…]
Here, the constraint  (9b) defines the binary indicator $a_{u,v}$, which specifies whether an ED is associated with a GW. Constraint  (9c) ensures that each ED can access at most one channel. Constraint  (9d) limits the number of EDs $v$ that can be associated with a single GW $u$ to at most $\Lambda_{max}$. Furthermore, constraint  (9e) defines the available set of SFs. Constraint  (9f) restricts the transmit power of the ED to be selected from a predefined discrete set. Constraint  (9g) ensures that the bandwidth allocated to each ED $v$ is chosen from a discrete set W. Finally, constraint  (9h) enforces that the received SNR must not fall below the threshold $\text{SNR}_{threshold}$, as shown in Table I, to ensure correct detection of LoRa EDs adopting a specific SF [6].
The formulated optimization problem (9) is NP-hard due to its combination of binary and stochastic constraints, along with the selection of discrete variables for SF, TP, and W, leading to a non-convex objective function that is challenging for conventional optimization methods. As established in [11], no efficient polynomial-time solution exists for this problem. Consequently, traditional RL methods, which require storing all MDP tuples in a table, are computationally intensive and unsuitable for multi-agent scenarios; thus, we design POMDP method and leverage the MAPPO-based algorithm to tackle the formulated problem in Eq.(9).
TABLE I: SINR Thresholds with $W=125$ kHz [6]
SF 7 8 9 10 11 12
SNRthreshold (dB) -7.5 -10 -12.5 -15 -17.5 -20
IV Multi-Agent PPO Algorithm
MAPPO is a promising framework that builds on the centralized training, decentralized execution (CTDE) scheme and extends the PPO to the multi-agent system [12, 13]. It focuses on the actor-critic architecture, which consists of two components: the actor, who is in charge of making decisions, and the critic, who analyses those actions using a value function. In a CTDE architecture shown in Fig. 2, the critic network learns a centralized value function that includes knowledge of all agents’ activities, whilst each agent uses the actor-network to determine its policy based solely on local observations. Furthermore, due to the limited communication range and the maximum number of EDs each flying LoRa GW can associate with at any given time, the global state of the system cannot be fully observed by a single agent. Consequently, we model the problem using a POMDP. In this framework, each agent only has partial information about the environment, leading to uncertainty in decision-making. Formally, the POMDP is defined by the tuple $\left\langle\mathcal{U},\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma,\mathcal{O}\right\rangle$.
Hence, the proposed framework consists of the following elements:
1) Agents $\mathcal{U}$: the set of flying LoRa GWs.
2) Observations $\mathcal{O}$: Due to partial observability, each flying LoRa GW $u$ at time $t$ perceives only a subset of the global state $\mathbf{\mathcal{S}}(t)$. Hence, the local observation $o_{u}(t)\in\mathcal{O}_{u}\subset\mathcal{S}(t)$ for agent $u$ is denoted as $o_{u}(t)=\left\{\psi(t),p(t),\Re(t),assEDs_{pos}(t),GWs_{pos}(t)\right\}$. Here, $assEDs_{pos}(t)$ represents the positions of the EDs currently associated with GW $u$, and $GWs_{pos}(t)$ are the positions of neighboring GWs within communication range. This partial observation mechanism ensures that each GW makes decisions based on its local view of the environment.
3) Actions $\mathcal{A}$: Each agent $u\in\mathcal{U}$ selects an action $a_{u}(t)\in\mathcal{A}_{u}$ at time $t$. The action comprises the selection of communication parameters including spreading factor, transmission power, and bandwidth. This can be represented as $a_{u}(t)=\left\{\psi_{u}(t),p_{u}(t),w_{u}(t)\right\}$, where $\psi_{u}(t)$, $p_{u}(t)$, and $w_{u}(t)$ denote the SF, TP, and W chosen by agent $u$ at time $t$, respectively.
4) Reward Function $\mathcal{R}$: The reward for each flying LoRa GW $u$ at time step $t$ is defined as:
$r_{u}(t)=\frac{\sum\limits_{v\in\mathcal{V}}\Re_{u,v}(t)\cdot a_{u,v}(t)}{\sum\limits_{v\in\mathcal{V}}P_{u,v}(t)\cdot a_{u,v}(t)+P_{c}}.$ (10)
Note that the reward in Eq.(10) is allocated only if the SNR constraint (9h) is satisfied; otherwise, the reward is zero. Therefore, the cumulative reward over the entire time horizon $T$ and all GWs $\mathcal{U}$ is:
$\mathcal{R}=\sum_{t=1}^{T}\sum_{u=1}^{U}\gamma^{t}r_{u}(t).$ (11)
As shown in Fig. 2, our system leverages MAPPO, which is based on the CTDE framework. During the training phase, MAPPO alternates between optimizing the actor and critic networks until stable convergence is achieved. Specifically, the flying GWs locally update the policy actor using the PPO method. At each training step, given the state of a GW, the actor-network is trained, and action is sampled according to the policy function $\pi_{\theta_{u}}(a_{u}(t),o_{u}(t))$. Once the joint action is executed, the corresponding reward is observed. Subsequently, the global state vector, representing the collective states of all flying GWs, is passed to the critic network. The critic network is trained by minimizing a predefined loss function, which helps evaluate future rewards for the policy and improves the overall strategy over time.
Figure 2: Centralized Critic and Decentralized Actor for MAPPO training.
V Performance Evaluation
V-A Simulation setup
[…]
In our simulations, the LoS and NLoS path loss coefficients are set as $\eta_{\text{LoS}}=0.1$ and $\eta_{\text{NLoS}}=21$, respectively. Additionally, we assume the path loss exponent factor $\lambda=0.43$, and the environment-dependent constant $\alpha=4.88$ [9]. Furthermore, we consider a carrier frequency of $f=868\,\text{MHz}$, the speed of light $c=3\times 10^{8}m/s$, and the noise power is set to $\sigma^{2}=10^{-15}W$. Furthermore, the agents can dynamically select configurations from three adjustable parameters. Particularly, SF is chosen from $\{7,8,9,10,11,12\}$, TP is selected from $\{2,5,8,11,14\}\,\text{dBm}$, and bandwidth $W$ configured from $\{125,250,500\}\,\text{kHz}$.
To train our agents, we employ a neural network architecture based on a gated recurrent unit (GRU) with 128 hidden units, a learning rate of $\alpha=3\times 10^{-4}$ for both actor and critic networks, and a soft target update coefficient of $0.01$. The training protocol uses a PPO clip range of $\epsilon=0.2$, a discount factor of $\gamma=0.99$, and the Adam optimizer.
As benchmarks, we evaluate our approach using two distinct schemes: the popular DRL-based algorithms and the conventional allocation schemes. For the former, we used MAA2C [14], COMA [15], and QMIX [16]. The MAA2C algorithm extends A2C to multi-agent scenarios by incorporating a centralized critic that learns the joint value function, while each agent maintains its own actor to learn individual policies. COMA is an actor-critic method where a centralized critic estimates the Q-function, and decentralized actors optimize their policies. QMIX is a deep RL method that promotes cooperative behavior by decomposing the value function in a way that preserves monotonicity. For the latter, we used a genetic-based algorithm and a distance-based allocation. Specifically, the distance-based approach assigns higher SFs and TP to more distant EDs. On the other hand, the genetic algorithm assigns different SFs, TP, and W to EDs to satisfy the SNR constraint in (9h). This algorithm is configured with a population size of 50, an elite size of 5, a mutation rate of 0.1, and 300 generations.
V-B Simulation Results
In Fig. 3(a), we present the association between EDs and GWs using a one-to-many matching association scheme, with a utility function designed to maximize the SNR. As illustrated in the figure, the quota constraint for upper-layer nodes (GWs) is respected, and all LoRa EDs are associated with at least one GW in such a way that the overall SNR of the network is maximized.
In Fig. 3(b), we depict the energy efficiency of the studied framework against environment timesteps for varying numbers of LoRa EDs, while keeping the number of GWs fixed at two. The figure shows that the agents experience different network topologies involving dynamic ED associations, mobility, and communication parameters during training. It can be observed that energy efficiency generally increases as the training steps evolve, with most network topologies reaching convergence around $1\times 10^{6}$ timesteps. The agents successfully learn optimal policies for allocating SF, TP, and W, considering both energy efficiency and the A2G wireless link. It is important to highlight that networks with lower ED density achieve higher energy efficiency, whereas those with higher ED density experience reduced energy efficiency. This outcome is logical given the increased complexity and higher overall power consumption introduced by denser network conditions.
In Fig. 3(c), we plot the energy efficiency against training timesteps for various network topologies. For example, we fixed 60 ED positions and varying numbers of GWs. As expected, configurations with a greater number of GWs converge faster and achieve higher energy efficiency, whereas topologies with fewer GWs converge more slowly. This trend is explained by the increased flexibility for EDs to associate with the nearest GW when more GWs are present. Such associations minimize transmission power required during data collection, simultaneously adhering to quota constraints and maximizing the utility function.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Refer to caption Refer to caption Refer to caption
(d) (e) (f)
Figure 3: (a) 3D plane of LoRa ED-GW association, (b) energy efficiency comparison with varying active EDs, (c) EE comparison with varying active GWs, (d) Cumulative power consumption with varying active EDs, (e) Training reward with different MARL benchmarking, (f) energy efficiency comparison with MARL and Non-MARL benchmarking.
Fig. 3(d) further evaluates the power consumption of the studied approach with 2 fixed GWs and a varying number of EDs. As illustrated, network topologies with higher ED densities initially consumed larger amounts of power, slightly above 1 watt. The increased initial consumption is logical due to the greater number of devices requiring communication. However, as training progresses, the agents effectively learn new policies that minimize power usage, thereby significantly improving the overall energy efficiency of the system. In contrast, topologies with fewer EDs consumed only a small amount of power from the start. Specifically, the zoomed-in plot shown in Fig. 3(d) clearly demonstrates the achievement of more realistic and efficient policies across all tested LoRa network topologies.
In Figs. 3(e), we compare the performance of our studied approach against other popular MARL algorithms. First, the cumulative reward functions are plotted against environment timesteps for DRL-based algorithms in Fig.3(e), as this metric applies exclusively to DRL-based approaches. The figure clearly indicates that our proposed method consistently outperforms state-of-the-art MARL algorithms, demonstrating both faster convergence and greater stability. Algorithms such as QMIX fail to converge to optimal policies due to the highly dynamic nature of the environment studied. Additionally, MAA2C exhibits a steadily increasing reward trend but does not converge to an optimal policy within the observed training period.
Furthermore, Fig. 3(f) illustrates a comparison of energy efficiency values against the number of EDs for DRL-based algorithms and other conventional allocation schemes. Our method consistently surpasses both the DRL-based approaches (MAA2C, COMA, and QMIX) and conventional allocation schemes (genetic algorithm and distance-based allocation). The significantly poorer performance of conventional schemes, particularly as the number of EDs increases, is logically consistent with their limited capability in handling complex, high-dimensional, and dynamically changing environments, as exemplified by our proposed LoRa system.
VI Conclusion
In this paper, we considered the joint optimization of assigning spreading factor, transmission power, bandwidth, and ED associations between multiple EDs and GWs while considering resource constraints and the A2G propagation to optimize the global system energy efficiency in a multi-flying LoRa network. We showed that the resulting sequential decision-making problem can be modeled as a POMDP, and we proposed a model-free RL algorithm that leverages a novel MAPPO scheme. We used simulation to show that our proposed approach can learn a good policy approximation by optimizing the global system EE in LoRa networks. Interesting directions of future work include integrating trajectory optimization for multiple flying LoRa gateways for efficient resource management. Our approach could be extended by exploring the Age of Information (AoI) scheme to enhance the freshness of data collection.
VII Acknowledgement
This work was sponsored by the Junior Faculty Development program under the UM6P-EPFL Excellence in Africa Initiative.